# T3: A Novel Zero-shot Transfer Learning Framework Iteratively Training on an Assistant Task for a Target Task

### v1

@@ A.3 Prompt for Summarization Generation and Experience Updating in Training Process
You are one helpful text assistant. Please help us to do the following:
1. Based on input text, provided questions, generated QA pairs and summary generation experience, generate summary corresponding answers to original text that can reflect all the subjects and all the key points of input article in proper length.
2. Update the summary generation experience of how to generate better summary with the help of provided questions first by the generated summary and all provided questions (the generated and the original) and then by what you learn when generating questions in question generation experience.
@@
THE GENERATED SUMMARY SHOULD HAS HIGHER ROUGE AND BLEU SCORES ALIGNED WITH THE ORIGINAL TEXT THAN THE DIRECTED GENERATED ONES. The summary should include essential terms, phrases, and sentences that closely mirror the original text to ensure high precision and recall. Notice Aim to maximize the overlap with the reference summary by closely following the wording, sentence structure, and style of the original text. Ensure that the summary is concise but comprehensive, avoiding unnecessary details or deviations. You need also consider the fluency of summary.
You need to consider all the key points to make a good summary and the learning experience should be general.(for example, like does not lose the detail in the story and pay attention to the connection of sentences, listing some point to improve the accuracy and the fluency of the story, YOU STILL NEED TO LIST OTHER rules THAT MAY IMPORTANT,Every point should as detailed as possible).
You can also think some methods to ensure a better summary, such as voting.
Answer in one line and point by point. DO NOT EXCEED 600 WORDS.’ ’The format of the output is as follows:
@@ A.4 Prompt in Test Process on Summarization Dataset
You are a helpful text assistant. Please help us to do the following:
Based on the input text, with the help of QA generation rules and summary rules, here are the steps to do the summary:
First, generate QA pairs based on the generation rules.
Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text in proper length.
@@ A.5 Prompt in Test Process on QA Dataset
You are a helpful text assistant. Please help us to do the following:
Based on the input text and the questions given, with the help of QA generation rules and summary rules, here are the steps to do the summary:
First, generate QA pairs based on the generation rules.
Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text in proper length.
@@ A.6 Summary Experience for News-style Datasets
1. Ensure all key points such as the formation, evolution, and current status of Boko Haram, and the contrasting views on security are included.
2. Maintain original wording and sentence structure for precision and recall.
@@
7. Verify the summary reflects the original text accurately without speculative information.
8. Address distinct key points or subjects in each question for comprehensive coverage.
9. Compare generated questions with original ones to ensure all critical aspects are covered.
10. Ensure answers are directly supported by the text, avoiding speculative or inferred information.
@@ A.7 Summary Experience for Narratives-style Datasets
1. Ensure all key points and characters from the text are covered to avoid missing crucial details,use Details to Enhance Emotional Impact such as Incorporate symbolic details.
@@
4. Keep the summary concise but comprehensive, avoiding unnecessary details while ensuring all significant events are included.
@@
10. Aim to maximize ROUGE and BLEU scores by closely following the wording, sentence structure, and style of the original text.
11. Highlight the emotional tones and themes present in the original text. This adds depth to the summary and helps convey the underlying messages or morals of the story. It’s important not just to recount events but also to capture the feelings and ideas that these events evoke.
12. Ensure that no important details are lost that contribute to the storyś moral or lesson. These details are often critical to understanding the overall message of the text, so they should be preserved to maintain the integrity of the original story.
13. Consider using a voting method to select the best summary from multiple versions, such as generate 3 versions and find the best one.
@@ A.8 QA Generation Experience for News-style Datasets
5. Answers should be directly supported by the text and avoid speculative or inferred information. This approach ensures comprehensive and accurate QA pairs.
@@ A.9 QA Generation Experience for Narratives-style Datasets
1.Ensure each question addresses a unique key point or detail to avoid redundancy. Cover all significant events, characters, and details for comprehensive coverage.
@@
4. Review the text multiple times to ensure no key points are missed and questions flow logically.
@@
6. Consider the context and sequence of events to create questions reflecting the narrativeś progression.
@@
9. Aim to create questions that encourage critical thinking and deeper engagement with the text.’

### v3

@@ A.3 Prompt for Summarization Generation and Experience Updating in Training Process
You are a helpful text assistant. Please help us with the following:
1. Based on the input text, provided questions, generated QA pairs, and summary generation experience, generate a summary that corresponds to the original text, reflecting all subjects and key points of the input article at an appropriate length.
2. Update the summary generation experience with insights on how to create a better summary by first using the provided questions, then the generated summary, and all provided questions (both generated and original), and finally incorporating what you’ve learned during question generation in the question generation experience.
@@
THE GENERATED SUMMARY SHOULD HAVE HIGHER PERFORMANCE SCORES COMPARED TO THE DIRECTLY GENERATED ONES. The summary should include essential terms, phrases, and sentences that closely mirror the original text. Aim to maximize the overlap with the reference summary by closely following the wording, sentence structure, and style of the original text. Ensure that the summary is concise but comprehensive, avoiding unnecessary details or deviations. Also, consider the fluency of the summary.
Consider all key points to make a good summary, and make the learning experience general (e.g., avoid losing details in the story, maintain sentence coherence, and list some points to improve accuracy and fluency). Additionally, you should list other rules that might be important, with each point as detailed as possible.
You may also consider methods to ensure a better summary, such as voting.
Answer in one line and point by point. DO NOT EXCEED 600 WORDS.
The format of the output is as follows:
@@ A.4 Prompt in Test Process on Summarization Dataset
You are a helpful text assistant. Please help us with the following:
Based on the input text, and with the help of QA generation rules and summary rules, here are the steps to create the summary:
First, generate QA pairs based on the generation rules. Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text at the appropriate length.
@@ A.5 Prompt in Test Process on QA Dataset
You are a helpful text assistant. Please help us with the following:
Based on the input text and the questions given, and with the help of QA generation rules and summary rules, here are the steps to create the summary:
First, generate QA pairs based on the generation rules. Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text at the appropriate length.
@@ A.6 Summary Experience for News-style Datasets
1. Ensure all key points, such as the formation, evolution, and current status of Boko Haram, as well as contrasting views on security, are included.
2. Maintain the original wording and sentence structure for precision and recall.
@@
7. Verify that the summary reflects the original text accurately without speculative information.
8. Address distinct key points or subjects in the summary for comprehensive coverage.
9. Compare generated summary with original text to ensure all critical aspects are covered.
10. Ensure summary are directly supported by the text, avoiding speculative or inferred information.
@@ A.7 Summary Experience for Narratives-style Datasets
1. Ensure all key points and characters from the text are covered to avoid missing crucial details; use details to enhance emotional impact, such as incorporating symbolic details.
@@
4. Keep the summary concise yet comprehensive, avoiding unnecessary details while ensuring all significant events are included.
@@
10. Closely follow the wording, sentence structure, and style of the original text.
11. Highlight the emotional tones and themes present in the original text. This adds depth to the summary and helps convey the underlying messages or morals of the story. It’s important not just to recount events but also to capture the feelings and ideas these events evoke.
12. Ensure that no important details contributing to the story’s moral or lesson are lost. These details are often critical to understanding the overall message of the text, so they should be preserved to maintain the integrity of the original story.
13. Consider using a voting method to select the best summary from multiple versions, such as generating three versions and choosing the best one.
@@ A.8 QA Generation Experience for News-style Datasets
5. Answers should be directly supported by the text and should avoid speculative or inferred information. This approach ensures comprehensive and accurate QA pairs.
@@ A.9 QA Generation Experience for Narratives-style Datasets
1. Ensure each question addresses a unique key point or detail to avoid redundancy. Cover all significant events, characters, and details for comprehensive coverage.
@@
4. Review the text multiple times to ensure no key points are missed and that questions flow logically.
@@
6. Consider the context and sequence of events to create questions reflecting the narrative’s progression.
@@
9. Aim to create questions that encourage critical thinking and deeper engagement with the text.
A.10 Step-by-step example
This section provides a detailed use case description, which helps readers understand the operation flow of each step more intuitively by analyzing the application of T3 step by step.
A.10.1 Original Text
In this example, we apply one sample from the dataset NarraSum, a story containing more than 1700 words, as our summarization target.
"The opening scene depicts Anita N̈eedyL̈esnicky (Amanda Seyfriend), a violent inmate in a correctional facility. She is notorious for being a ’kicker’, and is now reflecting during lockdown on past events. Needy is a nerdy, substantially unpopular, p̈lain Janeẗype student from Devil’s Kettle, Minnesota, who has been best friends with Jennifer Check (Megan Fox), a beautiful, popular cheerleader, since their childhood. …… The end credits show Low Shoulder checking into a hotel post-concert and candidly indulging in drugs and alcohol in their hotel room. Eventually, the band is savagely murdered by a blurred figure. We then see police crime scene photos of the band members’ corpses strewn around the hotel room, Nikolai having been stabbed in the chest with the sacrificial knife. The final image of the movie shows security camera footage of an evil-looking Needy walking down the hotel hallway past a gathering of female band groupies."
A.10.2 Generate QA Pairs
Using the sample and the QA generation experience in A.9 as input, the baseline model creates several question-answer pairs to capture key plot points and character details. Here are some examples:
Question: "What is Needy’s situation at the beginning of the movie?"
Answer: "She is an inmate in a correctional facility, known for her violent behavior as a ’kicker’."
Question: "Who are the main characters and their relationship?"
Answer: "Needy Lesnicky, an unpopular student, and Jennifer Check, a popular cheerleader, have been best friends since childhood.
Question: "What happens to the band Low Shoulder at the end?"
Answer: "The band is murdered by a figure, later revealed as Needy, after a concert while indulging in drugs and alcohol.
A.10.3 Final Summary
Using the generated QA pairs and the summary experience in A.7 as input, the baseline model generates a summary of the story as follows:
The story revolves around Anita N̈eedyL̈esnicky, a violent inmate reflecting on past events. Needy, a nerdy student from Devil’s Kettle, Minnesota, was best friends with Jennifer Check, a popular cheerleader. One night, they attend a concert by indie rock band Low Shoulder at a local bar. The bar catches fire, and Jennifer leaves with the band despite Needy’s protests. Later, Jennifer appears at Needy’s house, injured and behaving strangely. The next day, Jennifer acts normal at school but soon begins killing and feeding on male students. Needy becomes suspicious and investigates, discovering that Low Shoulder had attempted to sacrifice Jennifer in a Satanic ritual for fame and success. However, because Jennifer wasn’t a virgin, the ritual backfired, turning her into a succubus.Needy tries to warn her boyfriend Chip about Jennifer’s true nature, but he doesn’t believe her. At the Spring Formal, Jennifer attacks Chip, leading to a confrontation between Needy and Jennifer. Chip dies in Needy’s arms after trying to save her. Consumed by revenge, Needy breaks into Jennifer’s house and kills her. Needy is institutionalized for Jennifer’s murder but develops supernatural abilities from being bitten during their fight. She escapes the institution and sets out to seek revenge on Low Shoulder. The story concludes with security footage showing Needy walking down a hotel hallway, implying she has murdered the band members. The narrative explores themes of friendship, betrayal, and the supernatural, while also critiquing the music industry and teenage social dynamics. It blends elements of horror, dark comedy, and coming-of-age drama to create a unique and compelling story.
This final summary keeps the plot concise while covering essential character arcs, the supernatural element, and the revenge ending. As shown in this example, by using QA to capture plot intricacies, the proposed T3 framework ensures that the final summary is both comprehensive and aligned with the main themes of the original content.
A.11 Ablation study
Table 2: Ablation study results on BBC summary, NarraSum, NLQuAD and FairytaleQA
Models Strategy ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
GPT-3.5-turbo NLQuAD
Full T3 22.21 8.30 20.69 0.27 47.80
w/o summary 31.68 15.57 30.24 5.31 46.36
w/o qa 26.01 9.59 23.91 0.83 45.35
BBC Summary
Full T3 37.88 16.95 34.78 8.58 47.80
w/o summary 40.62 22.74 37.05 12.66 55.49
w/o qa 36.00 15.07 32.39 8.14 46.08
FairyTaleQA
Full T3 12.23 3.07 10.74 0.04 97.37
w/o summary 15.11 3.56 13.81 0.12 98.34
w/o qa 12.45 2.83 10.63 0.03 97.27
NarraSum
Full T3 25.28 5.88 23.00 2.15 95.91
w/o summary 26.40 6.34 23.67 2.27 94.86
w/o qa 23.28 5.14 20.73 1.98 93.69
GPT-4-turbo NLQuAD
Full T3 23.66 7.36 21.62 0.36 49.16
w/o summary 27.78 9.83 27.31 0.64 46.00
w/o qa 29.78 9.03 25.61 0.12 47.22
BBC Summary
Full T3 36.83 14.14 33.27 7.69 69.01
w/o summary 37.49 16.46 35.97 8.72 63.56
w/o qa 36.45 15.26 34.07 7.01 62.50
FairyTaleQA
Full T3 14.14 3.36 12.84 0.10 97.67
w/o summary 16.14 3.66 14.44 0.19 97.49
w/o qa 15.52 3.18 13.14 0.11 96.52
GPT-4-turbo NarraSum
Full T3 27.27 7.42 24.45 3.60 97.44
w/o summary 28.12 5.99 24.90 2.36 95.32
w/o qa 27.90 5.96 24.02 1.98 94.28
GPT-4o NLQuAD
Full T3 31.00 13.03 29.27 1.42 63.95
w/o summary 36.28 15.29 34.02 3.09 77.85
w/o qa 41.16 20.29 38.99 5.31 65.23
BBC Summary
Full T3 41.71 21.45 39.44 10.94 70.12
w/o summary 40.17 19.98 38.35 9.13 68.56
w/o qa 42.13 20.40 39.02 11.06 66.37
FairyTaleQA
Full T3 14.86 3.72 13.78 0.07 98.92
w/o summary 23.80 9.47 24.14 0.23 97.35
w/o qa 20.60 6.32 19.32 0.08 96.83
NarraSum
Full T3 28.86 8.58 25.66 2.55 99.69
w/o summary 27.44 7.28 24.83 2.45 98.50
w/o qa 29.03 7.22 26.04 2.71 97.42
Gemini-1.0-pro NLQuAD
Full T3 28.11 11.05 26.48 1.25 71.64
w/o summary 25.50 8.99 23.96 0.42 66.43
w/o qa 25.21 8.78 23.45 0.45 68.09
BBC Summary
Full T3 33.58 13.92 31.17 7.22 55.61
w/o summary 31.67 11.98 30.06 6.38 45.87
w/o qa 31.16 10.31 29.85 5.86 42.15
FairyTaleQA
Full T3 13.43 2.43 13.39 0.14 94.24
w/o summary 13.97 3.13 13.07 0.05 96.48
w/o qa 13.02 2.12 12.18 0.01 95.07
NarraSum
Full T3 28.25 5.74 26.36 2.09 91.22
w/o summary 27.63 5.91 24.59 1.67 93.70
w/o qa 26.73 5.47 24.06 2.08 92.57
Gemini-1.5-pro NLQuAD
Full T3 24.19 7.97 22.44 0.47 65.09
w/o summary 25.51 8.00 24.48 0.35 68.40
w/o qa 26.98 8.08 24.93 0.42 67.36
BBC Summary
Full T3 35.10 12.58 31.69 6.47 60.49
w/o summary 37.16 14.76 33.21 7.03 68.21
w/o qa 37.14 15.77 34.70 9.02 65.94
FairyTaleQA
Full T3 17.56 4.34 15.46 0.13 98.53
w/o summary 14.63 3.24 13.94 0.04 97.74
w/o qa 14.50 3.20 13.50 0.02 97.20
Gemini-1.5-pro NarraSum
Full T3 26.95 5.39 25.32 1.99 93.57
w/o summary 25.23 5.97 24.37 2.19 93.04
w/o qa 26.36 6.00 26.04 2.00 92.61
Claude-3.0-sonnet NLQuAD
Full T3 33.72 13.27 31.22 2.37 61.69
w/o summary 35.42 14.30 32.81 2.51 68.94
w/o qa 33.37 12.20 30.82 3.17 63.14
BBC Summary
Full T3 40.90 18.08 37.39 10.74 50.43
w/o summary 40.94 18.85 36.98 10.06 52.32
w/o qa 40.27 17.63 36.93 10.09 48.74
FairyTaleQA
Full T3 18.91 4.90 17.35 0.33 98.92
w/o summary 18.24 4.35 16.83 0.24 99.36
w/o qa 15.21 3.82 14.58 0.08 97.54
NarraSum
Full T3 25.87 5.47 22.51 2.05 95.61
w/o summary 28.34 6.57 26.05 2.47 97.59
w/o qa 27.52 6.09 25.21 2.02 94.83
Claude-3.5-sonnet NLQuAD
Full T3 40.17 18.74 37.92 5.00 82.76
w/o summary 36.79 16.71 34.45 4.09 75.03
w/o qa 38.08 15.83 35.80 4.02 72.94
BBC Summary
Full T3 43.54 21.13 40.36 13.00 65.83
w/o summary 40.26 16.83 35.99 13.31 58.32
w/o qa 41.55 20.72 40.05 31.90 58.57
FairyTaleQA
Full T3 20.96 6.07 19.52 0.47 99.97
w/o summary 19.94 5.84 19.41 0.53 98.53
w/o qa 19.86 5.36 18.18 0.31 97.98
NarraSum
Full T3 29.78 7.79 27.23 3.23 98.14
w/o summary 29.63 7.60 27.86 2.96 96.43
w/o qa 29.38 7.43 26.53 2.41 97.61
A.12 Significance Analysis
In this analysis, we used the t-test to evaluate the significance of the observed results. The null hypothesis $H_{0}$ assumes that there is no significant difference between the two algorithms compared in our experiment. The test statistic $t$ is calculated as follows:
$t=\frac{\bar{x}_{1}-\bar{x}_{2}}{\sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}}$ (3)
[…]
We calculate the sample means $\bar{x}_{1},\bar{x}_{2}$ and compute the sample standard deviations $s_{1},s_{2}$. Then we plug values into Equation (3) to determine the test statistic $t$ and compare $t$-value with the critical value from the $t$-distribution at a given confidence level $\alpha=0.05$.
The p-value for each metric result between our direct summarization and summary with T3 are listed in Table 3. Here we sign the model with * if the p-values for the metric are less than 0.05.
Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
BBC summary
GPT-4o* 0** 0** 0** 0.0007 0.0004
GPT-4-turbo 0.0431 0.1714 0.0367 0.0836 0.0227
GPT-3.5-turbo* 0.0124 0.0388 0.0080 0.0082 0.0017
Gemini-1.5-pro* 0.0160 0.0089 0.0011 0.0157 0.0160
Gemini-1.0-pro* 0.0101 0.0191 0.0103 0.0495 0.0091
Claude-3.5-sonnet* 0.0042 0.0003 0.0043 0.0001 0.0003
Claude-3.0-sonnet 0.0795 0.1108 0.0292 0.0268 0.5309
NarraSum
GPT-4o* 0** 0.0009 0** 0.0407 0.0003
GPT-4-turbo* 0.0077 0.0114 0.0177 0.0418 0.3652
GPT-3.5-turbo* 0.0499 0.0431 0.0478 0.0456 0.0203
Gemini-1.5-pro* 0.0332 0.0074 0.0196 0.0482 0.0459
Gemini-1.0-pro 0.0198 0.0312 0.0422 0.0696 0.0086
Claude-3.5-sonnet 0.0182 0.0417 0.0152 0.1526 0.1040
Claude-3.0-sonnet* 0.0483 0.0099 0.0523 0.0154 0.0691
NLQuAD
GPT-4o* 0.0086 0.0084 0.0072 0.0087 0.0420
GPT-4-turbo* 0.0189 0.0417 0.0162 0.0488 0.0340
GPT-3.5-turbo* 0.0299 0.0441 0.0376 0.0276 0.0457
Gemini-1.5-pro* 0.0436 0.0468 0.0282 0.0261 0.0014
Gemini-1.0-pro* 0.0479 0.0420 0.0289 0.0181 0.0498
Claude-3.5-sonnet* 0.0061 0.0003 0.0034 0.0002 0.0212
Claude-3.0-sonnet* 0.0095 0.0403 0.0427 0.0193 0.0089
FairytaleQA
GPT-4o* 0.0083 0.0180 0.0100 0.0194 0.0124
GPT-4-turbo* 0.0073 0.0375 0.0185 0.0367 0.0356
GPT-3.5-turbo 0.0409 0.0524 0.0895 0.0451 0.1184
Gemini-1.5-pro* 0.0473 0.0296 0.0188 0.0298 0.0365
Gemini-1.0-pro* 0.0168 0.0151 0.0169 0.0355 0.0496
Claude-3.5-sonnet* 0.0204 0.0031 0.0239 0.0423 0.0225
Claude-3.0-sonnet* 0.0340 0.0176 0.0338 0.0147 0.0092
Table 3: Experimental Results of the p-value on BBC summary, NarraSum, NLQuAD and FairytaleQA datasets between the result of each LLM without and with T3. 0** means the p-value is less than $10^{-7}$